### v1

# Time-reversed information flow through a wormhole in scalar-tensor gravity

Abstract
This Letter aims to advance new properties of a class of Closed Timelike Curves recently discovered in scalar-tensor gravity [Universe 9, 467 (2023)]. Therein, it was shown that a wormhole acts a gateway between two time-mirrored worlds, where the two asymptotically flat sheets in the Kruskal-Szekeres diagram are glued antipodally along three directions—time $t$ and the polar and azimuth angles $(\theta,\,\varphi)$ of the 2-sphere—to form a wormhole throat. This contrasts with the standard embedding diagram which typically glues the sheets only along the $\theta$ and $\varphi$ directions. Crucially, due to the ‘gluing’ along the $t$ direction, the wormhole becomes a portal connecting the two spacetime sheets with opposite physical time flows, enabling the emergence of closed timelike loops which straddle the throat. We shall point out that this portal mathematically permits the possibility of backward propagation of information against time. This feature is ubiquitous for wormholes in scalar-tensor theories. In addition, we formulate the Feynman sum for transition amplitudes of microscopic particles in the proximity of a wormhole throat in which we account for timelike paths that experience time reversal.
Introduction—Closed timelike curves (CTCs) are mathematical constructs first emerging in the context of General Relativity via the pioneering works of van Stockum and Gödel [1, 2]. These curves represent paths through spacetime that loop back on themselves, allowing an object or observer to return to an event in its own past. Interests in CTCs were revived in the works of Thorne and Novikov in the context of the Morris-Thorne wormholes which are thought to occur in the presence of exotic matter (or in modified gravity) [3, 4, 6, 5, 7, 8]. Despite the problematic nature of CTCs, the mathematical validity sparked considerable interest among gravitational theorists [9, 16, 12, 11, 13, 14, 10, 15, 17, 18, 19]. The concept of CTCs serves as a “gedanken experiment”, compelling us to grapple with and reevaluate the foundational principles of general relativity and its modifications.
Soon after the works of Thorne and collaborators, Deutsch and Politzer attempted to investigate quantum effects associated with CTCs. They introduced the Deutsch-Politzer (DP) space—an abstract construct created through surgical manipulation on a two-dimensional Lorentzian spacetime, forming a cylinder-shaped ‘handle’ that bridges two spacelike regions [20, 21, 23, 22]. This process is visually depicted in Fig.  1, where Paths 3 and 4 represent CTCs. As an object moves along a CTC, proper time continues to cumulate (i.e., the time count registered by a clock carried along the path keeps increasing), but the object eventually returns to its original spacetime point. Whereas geometry pertains to local properties of spacetime, topology pertains to global ones; in crafting a ‘handle’, the DP surgery deliberately alters the topology of spacetime, allowing for CTCs to emerge.
While DP space, as a conceptual model, serves its purpose, it is preferable to derive alterations to spacetime topology from a concrete theory of gravity. In fact, in General Relativity, it was shown that through the Einstein field equations these spacetimes violate the Weak Energy Condition (WEC) [3]. Matter that violates the WEC is denoted exotic matter, and can spontaneously induce spacetime to acquire non-trivial topologies, such as the Morris-Thorne wormholes that facilitates CTCs [4, 8, 24, 5, 25].
In a series of recent papers [26, 27, 28], one of the authors extended Buchdahl’s work in pure $\mathcal{R}^{2}$ gravity [29], culminating in a static spherisymmetric vacuum solution. Notably, despite the absence of exotic matter, the higher-derivative terms in the theory formally violate the WEC, under certain conditions. This violation enables the construction of a two-way traversable wormhole, which we named the Morris-Thorne-Buchdahl (MTB) wormhole [30]. Importantly, the analytical solution has allowed us to construct a ($\zeta-$)Kruskal-Szekeres (KS) diagram, revealing a non-trivial topology. The $(\zeta-$)KS coordinates enable for an identification procedure that results in the wormhole throat bridging two asymptotically flat spacetime sheets with opposing flows of proper time, as we recently proposed in [31]. Physical processes occurring in the two spacetime sheets across the throat evolve in opposite time directions. In essence, the MTB wormhole throat serves as a “gateway to another time-reversed world”. More generally, we have found that this property is a universal feature for Brans-Dicke and scalar-tensor gravities [32]. It is important to note that embedding diagrams, a standard tool to visualize wormholes, areinadequate in exposing this feature, as we shall explain in a later section of this Letter.
Importantly, the construction of the CTC that we presented in [31] does not necessitate one wormhole mouth to move at high speed or be situated near a supermassive object to accumulate time dilation—a procedure commonly portrayed in the literature [4, 8]. The resulting CTC is of a novel type, offering intriguing avenues for exploration and analysis.
Figure 1: The Deutsch-Politzer surgery. In the left panel, a 2-dimensional Minkowski spacetime is incised along the two dotted segments. The banks of the cuts are then glued together to form a ‘handle’, as shown in the right panel. Paths 1 and 2 represent regular trajectories in Minkowski spacetime. Path 3 is self-crossing, resulting in a portion of it forming a closed timelike curve (CTC). Path 4 depicts an infinite loop.
Figure 2: Visualizing a Buchdahl-inspired spacetime. While the embedding diagram (lower panel) suggests ‘gluing’ at a fixed timeslice $T$, the $\zeta-$KS diagram (upper panel) makes the ‘gluing’ along the time direction, with the two thick red lines are identical antipodally. An MTB wormhole is formed by flipping Region (IIIa) upside down and gluing it with Region (Ia).Trajectory (ABCDEFA) forms a closed timelike loop.
The special Buchdahl-inspired metric—Consider the vacuum pure $\mathcal{R}^{2}$ field equation [26]
$\mathcal{R}\,\Bigl{(}\mathcal{R}_{\mu\nu}-\frac{1}{4}g_{\mu\nu}\mathcal{R}\Bigr{)}+g_{\mu\nu}\square\,\mathcal{R}-\nabla_{\mu}\nabla_{\nu}\mathcal{R}=0$ (1)
Among a variety of solutions [27, 32], the asymptotically flat Buchdahl-inspired solution to Eq. (1) expressed in the Kruskal-Szekeres (KS) coordinates is given by [26]
$ds^{2}=-4e^{-r^{*}(r)}\Bigl{(}1-\frac{r_{\text{s}}}{r}\Bigr{)}^{\frac{\tilde{k}+1}{\zeta}}\left(dT^{2}-dX^{2}\right)+\Bigl{(}1-\frac{r_{\text{s}}}{r}\Bigr{)}^{\frac{\zeta+\tilde{k}-1}{\zeta}}r^{2}d\Omega^{2}$ (2)
and
$T^{2}-X^{2}$ $=-e^{r^{*}(r)};\ \ \ \frac{T}{X}=\tanh\frac{t}{2r_{\text{s}}}$ (3)
which depend on a dimensionless (Buchdahl) parameter $\tilde{k}$, and $\zeta:=\left(1+3\,\tilde{k}^{2}\right)^{1/2}$, whereas the tortoise coordinate $r^{*}(r)$ involves a Gaussian hypergeometric function and is given in [26]. Note: per Eqs. (3), each pair of coordinates $(t,r)$ corresponds to two pairs of KS coordinates $(T,X)$ and $(-T,-X)$.
In Ref. [30] we showed that, for $\tilde{k}\in(-1,0)$, the WEC is formally violated. Simultaneously, the areal radius, $R(r)=r\,\Bigl{(}1-\frac{r_{\text{s}}}{r}\Bigr{)}^{\frac{\zeta+\tilde{k}-1}{2\,\zeta}}$, attains a (local) minimum at
$r_{*}=\frac{\zeta-\tilde{k}+1}{2\,\zeta}\,r_{\text{s}}$ (4)
which exceeds $r_{\text{s}}$ when $\tilde{k}\in(-1,0)$. In this scenario, the spacetime forms a two-way traversable wormhole, referred to as the Morris-Thorne-Buchdahl (MTB) wormhole, with its throat located at $r_{*}$ given in (4), to be briefly described below. An embedding diagram is shown in Fig.  2.
Restricting within the $(T,X)$ plane (viz. $d\theta=d\varphi=0$), the $\zeta-$KS diagram for the metric (2)–(3) is shown in Fig.  2. The diagram is self-explanatory [30]: (i) It is conformally Minkowski, with null geodesics being $dX=\pm dT$. (ii) A constant–$t$ contour corresponds to a straight line running through the origin of the $(T,X)$ plane. (iii) Regions (I) and (III) are mirror images of each other, upon flipping the sign of the KS coordinates, viz. $(T,X)\leftrightarrow(-T,-X)$. (iv) For $\tilde{k}\in(-1,0)$, the hyperbola $r=r_{*}$ (thick red lines) splits these regions into sub-regions (Ia), (Ib), (IIIa), (IIIb). The ‘outer’ sub-regions (Ia) and (IIIa) are asymptotically flat spacetime sheets which are devoid of physical singularities.
Figure 3: Penrose diagram illustrating an MTB wormhole with its two sheets presented as the left and right wedges. Constant-$r$ contours connecting $i^{-}$ with $i^{+}$ are shown. The wedges are glued along the thick red lines, corresponding to the throat at $r=r_{*}$, but with the left wedge and line flipped upside down, resulting in a non-trivial topology where the physical time direction flips across the throat. Path $(abc)$ is a CTC, with $(b^{\prime})$ mirroring $(b)$. The shaded area is not a part of the wormhole; it contains physical singularities at $r=0$ and $r=r_{\text{s}}$.
Identification of antipodal points and emergence of CTCs—Antipodal points play a crucial role in the $\zeta-$KS diagram. Specifically, the two thick red lines on the diagram, in Fig.  2, represent a wormhole throat, and the identification of antipodal points $(T,X)$ and $(-T,-X)$ on these lines signifies a single spacetime event occurring on the throat. It is important to note that this identification is limited strictly to points on the throat. Points that stay off the throat do not enjoy this privilege; they correspond to independent events occurring on separate spacetime sheets. (Note: Our procedure contrasts with historical maneuvers known as “elliptic/antipodal identification” [38, 36, 37, 33, 34, 35, 39]. In particular, when dealing with the regular KS diagram of a Schwarzschild spacetime, Rindler [33] assigned each antipodal pair to the same event unrestrictedly, whereas Popławski [34] applied such an assignment to antipodal points that lie on the Schwarzschild horizon.)
In Fig.  2, the trajectory $A\rightarrow B\equiv C\rightarrow D\rightarrow E\equiv F\rightarrow A$ constitutes a closed timelike loop. Note that Points B and C, as well as Points E and F, are antipodal on the throat and, therefore, identical. The segments between these points can freely take any form as long as they reside within the lightcones. One key feature is that while the segment $F\rightarrow A\rightarrow B$ progresses in an increasing order of the ‘time’ coordinate $t$, the other segment $C\rightarrow D\rightarrow E$ advances in a decreasing order of $t$. (Note: the proper time $\tau$ increases along $C\rightarrow D\rightarrow E$.) Essentially, the wormhole throat serves as a gateway to another time-reversed world by connecting two sheets with opposing physical time flows.
Akin to the DP space which incurs an alteration of spacetime topology, the $\zeta-$KS diagram acquires a non-trivial topology. However, rather than relying on abstract manipulation as in the DP space, the change in topology for our $\zeta-$KS diagram takes place spontaneously, driven by the inherent higher-derivative nature of the $\mathcal{R}^{2}$ theory. The pertinent sections of the Penrose diagram are shown in Fig.  3, with explanations provided in its caption. Owing to the antipodality, the left wedge must be flipped upside down to facilitate the gluing of the two thick red lines in forming an MTB wormhole.
It is worth noting that the opposite time flows for physical processes in Quadrants (I) and (III) have been known since the conception of the KS diagram in the 1960s. Yet, this fact has often been overlooked in wormhole studies, most likely because of the prevalent use of embedding diagrams, which are a standard technique for visualizing wormholes. Embedding diagrams obscure a fundamental feature, as they suggest ‘gluing’ the two sheets along the two angular directions, namely the polar and azimuth angles ($\theta$ and $\varphi$), at the minimal areal radius. However, in an MTB wormhole, the ‘gluing’ actually occurs in three directions – the angular ones ($\theta$ and $\varphi$) and the time direction ($t$). This only becomes evident with the explicit use of the $\zeta-$KS diagram. In Fig. 2, one of the thick red lines needs be flipped upside down, and then the two said lines get glued in their entirety, viz. along the time direction as well, besides $\theta$ and $\varphi$.
The portal connecting two time-reversed sheets is ubiquitous not only in pure $\mathcal{R}^{2}$ gravity but also in Brans-Dicke and scalar-tensor theories, as they share a similar KS diagram [32]. It is intuitive to expect that the existence of time-reversed portals is a common feature among wormholes at large. We will now present three corollaries arising from the existence of CTCs in the MTB wormhole.
Corollary #1: Perception of time reversal via light signal exchange—The upper panel of Fig.  4 exhibits the case. Consider a Traveler in the left sheet emitting light signals into the wormhole in the order $A$, $B$, $C$, and $D$; the light signals emerge into the right mouth and continue on their respective null paths. Note that on the throat, each antipodal pair correspond to a single event, viz. $A_{1}\equiv A_{2}$, $B_{1}\equiv B_{2}$, $C_{1}\equiv C_{2}$, and $D_{1}\equiv D_{2}$. An Observer residing in the right sheet receives light signals in a reversed order, $D^{\prime}$, $C^{\prime}$, $B^{\prime}$, then $A^{\prime}$, however. Therefore, from the Observer’s perspective, the Traveler appears to move backward in time. The perception is mutually symmetric: the Traveler would perceive the Observer as moving backward with respect to the Traveler.
Figure 4: Illustration of Corollaries #1 and #2. Upper panel: Traveler in the left sheet regularly sends light signals to Observer in the right sheet (Note: $A_{1}\equiv A_{2}$, $B_{1}\equiv B_{2}$ and so on.) The order of arrival for light signals is reversed. Lower panel: Traveler at $A$ sends out a light signal, then crosses through the throat. Later on, Traveler while passing by $A^{\prime}$ deflects the light signal back; the light signal arrives at $A^{\prime\prime}$ which precedes its emission event $A$. (Note: $A_{1}\equiv A_{2}$ and $A^{\prime}_{1}\equiv A^{\prime}_{2}$.)
Corollary #2: Backward flow of information against time—In the lower panel of Fig. 4, we depict another closely related effect. Consider a Traveler in the left sheet. At point $A$, he emits a light signal that enters the wormhole at $A_{1}$ and emerges at $A_{2}$ (note their antipodal nature). The Traveler then passes into the wormhole and escapes to the right sheet. Choosing a trajectory such that at point $A^{\prime}$ (where he crosses paths with the original light signal), he deflects the light signal back into the wormhole (e.g., using a mirror). The chosen path bends ‘outward’ to the right, delaying the reflection event. The deflected light signal at $A^{\prime}$ re-enters the throat at $A^{\prime}_{1}$ and re-emerges at $A^{\prime}_{2}$ (again, note their antipodal nature). Finally, the light signal arrives at point $A^{\prime\prime}$, preceding the emission event $A$.
Corollary #3: Bidirectional time flow on the throat—A Traveler sitting precisely on the throat could proceed either forward into the right sheet or backward into the left sheet. If choosing to reside on the throat, both directions of time are admissible for him.
Prevention of free will to act on foreknowledge?—There seems to be no inherent principle forbidding microscopic objects from “exploiting foreknowledge to their advantage”. Inanimate objects lack free will to alter history, such as murdering their younger selves or preemptively securing the winning lottery tickets. A timelike path involving ‘time travel’ for microscopic particles would pose no such threat and should be allowed to contribute in the Feynman sum of transition amplitudes for these particles, as we will formulate shortly. The implications of foreknowledge on macroscopic animate objects constitute a subject of ongoing debate and deliberation, nevertheless [9, 16, 12, 11, 6, 13, 14, 10, 7, 15, 17, 2, 18, 19]. We venture one view: macroscopic objects might undergo a strong decoherence process upon passing through the wormhole throat. If violent enough, the decoherence could erase, for instance, the mental state of an intrepid traveler crossing the throat. For example, in Corollary #2 above, memory erasure could derail the Traveler from carrying out his plan to deflect the light signal at $A^{\prime}$, thus thwarting its ‘early’ arrival at $A^{\prime\prime}$. This decoherence-based reasoning could be applicable to macroscopic objects in general, such as an unmanned spaceship pre-programmed to traverse to $A^{\prime}$ to deflect the light signal. The decoherence might erase its computer program and sabotage its planned itinerary.
[…]
$\tau_{AB}$ $=\int_{t_{A}}^{t_{B}}dt\,\Bigl{(}1-\frac{1}{r}\Bigr{)}^{\frac{\tilde{k}-1}{2\,\zeta}}\times$ (5)
$\ \ \ \Bigl{[}\Bigl{(}1-\frac{1}{r}\Bigr{)}^{\frac{2}{\zeta}}-\Bigl{(}\dot{r}^{2}+\Bigl{(}1-\frac{1}{r}\Bigr{)}r^{2}\bigl{(}\dot{\theta}^{2}+\sin^{2}\theta\,\dot{\varphi}^{2}\bigr{)}\Bigr{)}\Bigr{]}^{\frac{1}{2}}$
with the overdot indicating a derivative with respect to $t$. Denoting $z:=1-r_{\text{s}}/r\in[r_{*},1)$ with $z_{*}:=1-r_{\text{s}}/r_{*}$ and $r_{*}$ given in Eq.  (4) specifying the location of the throat, Eq.  (5) yields
$\tau_{AB}=\int\displaylimits_{t_{A}}^{t_{B}}\frac{dt\text{\,}z^{\frac{\tilde{k}-1}{2\,\zeta}}}{(1-z)^{2}}\Bigl{[}z^{\frac{2}{\zeta}}(1-z)^{4}-\dot{z}^{2}-z(1-z)^{2}\bigl{(}\dot{\theta}^{2}+\sin^{2}\theta\,\dot{\varphi}^{2}\bigr{)}\Bigr{]}^{\frac{1}{2}}$ (6)
The geodesics in the plane $\theta=\pi/2$, obtained by extremizing the functional (6), satisfies the following equations, with $l$ and $\mathcal{E}$ being two constants of motion:
$\dot{z}$ $=(1-z)^{2}\,z^{\frac{1}{\zeta}}\frac{1}{\mathcal{E}}\sqrt{\mathcal{E}^{2}-(1-z)^{2}z^{\frac{2-\zeta}{\zeta}}l^{2}-z^{\frac{\tilde{k}+1}{\zeta}}}$ (7)
$\dot{\varphi}$ $=(1-z)^{2}\,z^{\frac{2}{\zeta}-1}\frac{l}{\mathcal{E}}$ (8)
In the large $\mathcal{E}$ limit, Eqs.  (7)–(8) are integrable to yield $t-t_{0}=\frac{1}{1-1/\zeta}\,z^{1-1/\zeta}\,_{2}F_{1}\left(2,1-1/\zeta;2-1/\zeta;z\right)$ and $\varphi=\text{const}$, corresponding to a pure radial motion.
Figure 5: Different transition paths from Event A to Event B (both located in the right sheet of the wormhole). In the upper row, 3 paths exhibit no crossing into the left sheet, while another path involves a single crossing into the left sheet and subsequent return. The lower row presents 5 admissible arrangements for paths with two crossings into the left sheet and twice-returning. In all cases, traversal in the left sheet is depicted by a segment oriented against the direction of $t$, meaning that a time-reversed path entails a multi-valued function $z(t)$.
Consider now a pair of starting and end points $A$ and $B$ situated in the right sheet. Note that Eq.  (6) has been defined for a single-valued function $z(t)$ that connects $A$ with $B$ while staying above the line $z=z_{*}$. This scenario is depicted in the upper left plot in Fig.  5, showing 3 paths with no crossing into the left sheet (the boundary of which is the horizontal axis $t$, representing $z=z_{*}$).
When a path crosses the wormhole throat and ventures into the left sheet, the respective segment in the left sheet has a reversed time direction, however. This can conveniently be superimposed in the same $(t,z)$ plane except with an orientation against the direction of $t$. Other plots in Fig.  5 showcase the once-time-reversed and twice-time-reversed paths. Given the additive nature of proper time, it is straightforward to see, for example, that for the once-time-reversed path: $\tau_{AA_{1}A_{2}B}=\tau_{AA_{1}}+\tau_{A_{1}A_{2}}+\tau_{A_{2}B}$.
The transition amplitude of a particle of mass $m$, as usual, is given by a Feynman sum, $K(\vec{z}_{B},\vec{z}_{A};t_{B}-t_{A})\simeq\int\mathcal{D}\vec{z}(t)\,\exp\left(-i\,m\,\tau_{AB}(\vec{z}(t))\right)$. Note that in the absence of a wormhole throat, all paths that contribute to the Feynman sum must run forward in time; this means that the function $z(t)$ is single-valued. However, when facilitated by a wormhole, a time-reversed path would make $z(t)$ a multi-valued function, as is evident from Fig.  5. The contribution of time-reversed paths into the Feynman sum is thus a novel non-trivial aspect. Moreover, in the proximity of a throat, timelike paths with multiple time reversals could proliferate and significantly—if not fundamentally—alter the transition amplitude of free particles. This intuitive possibility underscores the importance of the decoherence process discussed in relation to safeguarding causality for macroscopic objects. We leave this arena open for future exploration.
[…]
$L:=-\frac{m\,z^{\frac{\tilde{k}-1}{2\,\zeta}}}{(1-z)^{2}}\left[z^{\frac{2}{\zeta}}(1-z)^{4}-\dot{z}^{2}-z(1-z)^{2}\,\dot{\varphi}^{2}\right]^{\frac{1}{2}}$ (9)
and, via the generalized momenta $p_{z}:=\frac{\delta L}{\delta\dot{z}}$ and $p_{\varphi}:=\frac{\delta L}{\delta\dot{\varphi}}$, the associated Hamiltonian is given by
$H$ $=p_{z}\,\dot{z}+p_{\varphi}\,\dot{\varphi}-L$ (10)
$=z^{\frac{1}{\zeta}}\sqrt{(1-z)^{4}\,p_{z}^{2}+z^{-1}(1-z)^{2}\,p_{\varphi}^{2}+m^{2}\,z^{\frac{\tilde{k}-1}{\zeta}}}$ (11)
with $p_{\varphi}$ and $H$ being two constants of motion.
Hamiltonians are known to be suitable for the Heisenberg picture when describing the evolution of free particles and for the interaction picture when building the Feynman rules for quantum field theories. Despite the apparent asymmetric treatment of the time direction (versus the spatial ones) by the Hamiltonian, relativistic covariance is duly restored in the final results [40]. In the context of an MTB wormhole, where the background metric is static and inert, the Hamiltonian could find utility. As a by-product, the form of the Hamiltonian in Eq.  (11) is also evocative of Dirac’s historical approach towards the (relativistic) Dirac equation in 1927. Starting from the energy-momentum relation $E^{2}=p^{2}+m^{2}$ and requiring the field equation for the electron to be first-order, Dirac conceived the spinors and the associated Clifford algebra for the Dirac gamma matrices in Minkowski spacetime. Notwithstanding the tetrad formalism, Expression  (11) suggests pathways for an alternative and convenient representation of Dirac’s spinor fields in the presence of an MTB wormhole. The Clifford algebra may be adapted to conform with the altered energy-momentum relation, obtained by ‘squaring’ Eq.  (11):
$z^{-\frac{2}{\zeta}}\,E^{2}=(1-z)^{4}\,p_{z}^{2}+z^{-1}(1-z)^{2}\,p_{\varphi}^{2}+m^{2}\,z^{\frac{\tilde{k}-1}{\zeta}}$ (12)
Lastly, the time-reversed information flows across a wormhole, as demonstrated in Corollary #2, could potentially impact the unitarity of the evolution operator and the time-ordered product in the interaction picture. These intriguing prospects warrant in-depth exploration.
Acknowledgments—H.K.N. thanks Tiberiu Harko, Mustapha Azreg-Aïnou, Luis A. Correa-Borbonet, and Nicholas Buchdahl. F.S.N.L. acknowledges support from the Fundaşão para a Ciência e a Tecnologia (FCT) Scientific Employment Stimulus contract with reference CEECINST/00032/2018, and funding from the research grants UIDB/04434/2020 and UIDP/04434/2020.